# Best-Worst Disaggregation: An approach to the preference disaggregation problem

### v1

@@ 3 Best-worst disaggregation method
where the $i$th row is the profile of the $i$th alternative, i.e., $\mathbf{x}_{i}$. Following the precepts of UTASTAR, we identify a subset of alternatives and we call $R\subseteq M$ its index set. Then, according to the principles of BWM, we identify the indices of the best and the worst alternatives in the reference set and we label them with the letters $B$ and $W$, respectively. We then conduct a holistic comparison between the best alternative $\mathbf{x}_{B}$ and the other alternatives, as well as between all the alternatives and the worst alternative $\mathbf{x}_{W}$. Such comparisons can be collected into two sets as follows:
@@
Solving model (BWD), we can obtain the piecewise linear approximations of the attribute value functions and its optimum, $\xi^{*}$, is and indicator of the compatibility of the preferences of the expert with the most suitable underlying value function $V_{L}$. If the results are deemed reliable, then the obtained linear approximations can be used to rate the entire set of alternatives $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$. Coherently with the literature on preference disaggregation, we suggest that further analyses, especially on sensitivity, be carried out to test the robustness of the results.
@@ 3.1 A proposal to select the reference set
The choice of the reference set $R$ can be seen as a preprocessing phase, but this should not belittle its importance. While is existing disaggregation studies, the reference set is usually given, here, we consider some desiderata to propose a method to select the reference set:
@@
    For all the attributes, the selected alternatives should cover all the segments into which the attribute ranges were divided. This insures that there is preferential information on each segment of each attribute.
@@
    The cardinality of the reference set should be minimized to decrease the cognitive load. This is coherent with current parsimonious approaches to the elicitation of preferences (Abastante et al., , 2019; Corrente et al., 2024a, ).
@@
We propose a combinatorial optimization problem such that each segment is covered by at least $b\in\mathbb{N}_{+}$ alternatives and that in the selected set there is not any Pareto dominance relation between alternatives. That it, we can solve
@@ 3.2 Necessary relation from the BWD model
It could happen that optimization problem (BWD) yields infinitely many optimal solutions. For this reason we suggest an approach that tries to find the necessary ranking of the alternatives. That is, we propose an approach inspired by Robust Ordinal Regression (Corrente et al., , 2013). First, we call $\xi^{*}$ the optimal value of (BWD). For each pair of alternatives $(\mathbf{x}_{p},\mathbf{x}_{q})$, we maximize their value difference as follows:
@@ 3.3 Consistency analysis
In the previous section we recognized that the optimal value $\xi^{*}$ contains information on the discrepancy between the preferences expressed by the expert and the most suitable underlying model $V_{L}$. Hence, we can interpret $\xi^{*}$ as an index of compatibility. It is worth noting that compatibility should not be confused with the concept of (internal) consistency, which, in the case of the BWM collapses into the condition $a_{Bj}a_{jW}=a_{BW}$ for all $j\in N$. Consider, for instance, that the preferences of the expert can be perfectly consistent, i.e. rational, and yet violate the principle of Pareto dominance between alternatives at the same time. This latter violation implies that the preferences cannot be fully compatible with the underlying model.
@@
We further note that, in the realm of preference disaggregation, the evaluation of internal consistency is a specific feature of the BWD, and possibly an unicum in the field of preference disaggregation. Other models, such as UTA and UTASTAR, require, as input, preference orders which are, by definition, transitive. Furthermore, they provide ordinal information and thus any analysis of cardinal information loses its grip. We conclude, also in light of Figure 1, that the analysis of compatibility is not enough and it should be coupled with an analysis of the consistency of the preferences. Starting from the principle that the level of inconsistency of preferences should not be model-dependent, therefore, we adopt the Ordinal Consistency Ratio ($OR$) and the Cardinal Consistency Ratio ($CR$) proposed by Liang et al., (2020) to quantify deviations from these consistency conditions in the BWD method.
@@
3.4 Sensitivity analysis on the possible ranks of alternatives
In some decision-making contexts, it is often important to determine the maximum and minimum possible ranks that can be achieved by an alternative in the final ranking. This type of sensitivity analysis helps assess the stability of an alternative’s position. To determine these bounds, we solve two optimization problems for each alternative, always searching within the set of the optimal solutions of (BWD).
@@ 3.5 An interval-valued extension
subject to $-\xi+a_{Bi}^{-}V_{L}(\mathbf{x}_{j})\leq V_{L}(\mathbf{x}_{i})\leq a_{Bi}^{+}V_{L}(\mathbf{x}_{j})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
$-\xi+a_{iW}^{-}V_{L}(\mathbf{x}_{j})\leq V_{L}(\mathbf{x}_{i})\leq a_{iW}^{+}V_{L}(\mathbf{x}_{j})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
@@ 4 A case study in logistics performance evaluation
We tested the BWD on the data collected and used by Rezaei et al., (2018) to rank countries according to a composite indicator of logistic performance. In their study, data from the World Bank was used to assess the performance of the World countries with respect to the following criteria: Customs; Infrastructure; Services; Timeliness; Tracking and tracing; International shipment. For sake of simplicity and ease of interpretation of the results we only focused on European countries.
@@
One splitting point was allowed for each criteria set and was positioned midway in the interval $[\underline{x}_{j},\overline{x}_{j}]$. Next, we solved the set covering-inspired optimization problem (15)–(18) to find a mutually non-dominated minimum covering set. This set contained only three countries and was taken as a good initial suggestion. Given the possibility of asking for more comparisons we added two more alternatives to better explore the lower part of the ranking. The final reference set $R$ was defined as
@@
Using the Ordinal Consistency Ratio ($OR$) as formulated in eq. (20), we found $OR=0.2$, indicating a violation of ordinal consistency. The local ordinal consistency ratios ($OR_{i}$), shown in Table 2, help locate specific inconsistencies. In this case, the preferences provided by the expert on Latvia are not ordinal consistent with those expressed on Greece
@@
After identifying and locating inconsistencies in the preferences, we re-engaged with the expert to review and reconsider her judgments. Upon reflection, the expert acknowledged the inconsistencies and proceeded to modify the preferences using the consistency improvement process from the work of Liang et al., (2022). Using the same approach we derived the admissible ranges for $A^{BO}$ and $A^{OW}$, as shown in Table 3, and the visualization of the improving ranges, shown in Figure 2.
@@
With the help of the visualization of the ranges for improvement, the expert can easily adjust the inconsistent comparison. After consideration, the expert made some minor revision on the preferences. The modified sets $A^{BO}$ and $A^{OW}$ are reported in Table 4.
@@
The full analysis revealed that $\succ_{\mathcal{N}}$ is almost a total order relation. This is confirmed in Figure 3, showing very low variability in the set of rankings obtained through optimization problems (22) and (23). Only one swap is possible, between Greece and Slovenia. The results shown in Figure 3 seem to hint at a good precision of the outcome and further tests that we conducted, but we avoid to report, showed any level of incompatibility, i.e. $\xi^{*}>0$, very often leads to an extremely stable ranking.
@@
We also tested the I-BWD on the same case study, together with the analysis of dominance relations and maximum and minimum ranks. In this case, the expert information is summarized in Table 5. Note that the improving ranges shown in Table 3 can be used to guide experts to extend their original real-valued preferences to interval ones in a way that it may decrease their inconsistency.
@@
A full analysis of the maximum and minimum attainable ranks is presented in the right hand side of Figure 3 and confirms the previous intuition on the much greater variability of the attainable ranks. To quantify the imprecision due to the existence of multiple compatible rankings, we introduce the following heuristic index
@@
where each term of the sum is the “spread” of the possible ranks that the $i$th alternative can attain divided the maximum possible range, going from $1$, the best rank, to $m$, the lowest possible attainable rank. The scaling factor preceding the sum normalizes the quantity into the interval $[0,1]$. In this case of the I-BWD, we have $U=0.05938$, whereas, for the real-valued BWD it was $U=0.00135$.
@@ 5 Conclusions
The Best-Worst Disaggregation Method (BWD) offers an advancement in the field of Multi-Criteria Decision-Making (MCDM) by combining the principles of the Best-Worst Method (BWM) with the disaggregation approach. This method addresses the challenges inherent in traditional aggregation and disaggregation techniques by leveraging the strengths of both to create a robust and reliable decision-making framework. It is a valuable tool in scenarios where experts have a clear understanding of alternatives but may not have predefined value functions.
@@
Future research on BWD could focus on several key areas to enhance its applicability and effectiveness: extending BWD to accommodate group decision-making scenarios by developing techniques to aggregate individual preferences and address potential conflicts; adapting BWD for use in dynamic environments where decision criteria and alternatives may change over time; enhancing the computational efficiency of BWD to handle larger sets of criteria and alternatives by exploring advanced optimization techniques and parallel computing strategies; investigating methods to minimize biases in expert judgments, such as incorporating techniques from behavioral decision theory or utilizing machine learning or artificial intelligence to identify and correct inconsistencies; conducting practical case studies in various domains to validate the effectiveness of BWD and identify specific areas for improvement. Finally, we think that developing the BWD for group decision-making and for other problems such as sorting could be interesting areas for further development.

### v2

@@ 3 Best-worst disaggregation method
where the $i$th row is the profile of the $i$th alternative, i.e., $\mathbf{x}_{i}$. Following the precepts of UTASTAR, we identify a reference set of alternatives and we call $R$ its index set. Then, according to the principles of BWM, we identify the indices of the best and the worst alternatives in the reference set and we label them with the letters $B$ and $W$, respectively. Let us incidentally note that the freedom in the definition of the reference set $R$ allows the decision maker to choose it in such a way to simplify the choice of the best and the worst alternative within it. We then conduct a holistic comparison between the best alternative $\mathbf{x}_{B}$ and the other alternatives, as well as between all the alternatives and the worst alternative $\mathbf{x}_{W}$. Such comparisons can be collected into two sets as follows:
@@
Solving (BWD), we can obtain the piecewise linear approximations of the attribute value functions and its optimum, $\xi^{*}$, is an indicator of compatibility of the preferences of the expert with the most suitable underlying value function $V_{L}$. If the results are deemed reliable, then the linear approximations can be used to rate the entire set of alternatives $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$. Coherently with the literature on preference disaggregation, we suggest that further analyses, especially sensitivity, be carried out to test the robustness of the results.
It must be pointed out that Barbati et al., (2024) already suggested a formulation compatible with (BWD)—see eq. (14) in their paper—with the only differences that (i) they employ the Deck-of-Cards method to rate alternatives (ii) they consider the final rating obtained from the Deck-of-Cards method instead of a set of comparisons $A^{BO}\cup A^{OW}$. Hereafter we will focus on (BWD) and propose methods to study consistency, compatibility and extreme rankings. Furthermore, we stress that we employed a min-max approach to remain coherent with the original BWM but, as shown also by Barbati et al., (2024), the optimization problem can be easily modified to accommodate other formulations, e.g., the sum of the errors.
@@ 3.1 A proposal to select the reference set
The choice of the reference set $R$ can be seen as a preprocessing phase, but this should not diminish its importance. While in existing disaggregation studies, the reference set is usually given, here, we consider some desiderata to propose a method to select the reference set:
@@
    For all the attributes, the selected alternatives should cover all the segments into which the attribute ranges were divided. This ensures that there is preferential information on each segment of each attribute.
@@
    The cardinality of the reference set should be minimized to decrease the cognitive load. This is coherent with current parsimonious approaches to the elicitation of preferences (Abastante et al., , 2019; Corrente et al., , 2024).
@@
We propose a combinatorial optimization problem such that each segment is covered by at least $b\in\mathbb{N}_{+}$ alternatives and that in the selected set there is not any Pareto dominance relation between alternatives. That is, we can solve
@@
It is also worth mentioning that the purpose of the proposed approach here is to find a non-dominated set as a reference point, not to find the Pareto optimal set of the alternative set. That is, the remaining alternatives which are not selected as the reference set are not necessarily dominated alternatives. Lastly, a cautionary note: it may be increasingly difficult to solve this problem for a growing number of splitting points. In fact, the greater number of segments to be covered conflicts with the constraint that the chosen reference alternatives cannot dominate each other.
@@ 3.2 Necessary relation from the BWD model
It could happen that the optimization problem (BWD) yields infinitely many optimal solutions. For this reason, we suggest an approach that tries to find the necessary ranking of the alternatives. That is, we propose an approach inspired by Robust Ordinal Regression (Corrente et al., , 2013). First, we call $\xi^{*}$ the optimal value of (BWD). For each pair of alternatives $(\mathbf{x}_{p},\mathbf{x}_{q})$, we maximize their value difference as follows:
@@ 3.3 Consistency analysis
In the previous section we recognized that the optimal value $\xi^{*}$ contains information on the discrepancy between the preferences expressed by the expert and the most suitable underlying model $V_{L}$. Hence, we can interpret $\xi^{*}$ as an index of compatibility. It is worth noting that compatibility should not be confused with the concept of (internal) consistency, which, in the case of the BWM, collapses into the condition $a_{Bj}a_{jW}=a_{BW}$ for all $j\in N$. Consider, for instance, that the preferences of the expert can be perfectly consistent, i.e. rational, and yet violate the principle of Pareto dominance between alternatives at the same time. This latter violation implies that the preferences cannot be fully compatible with the underlying model.
@@
We further note that, in the realm of preference disaggregation, the evaluation of internal consistency is a specific feature of the BWD, and possibly a unicum in the field of preference disaggregation. Other models, such as UTA and UTASTAR, require, as input, preference orders which are, by definition, transitive. Furthermore, they provide ordinal information and thus any analysis of cardinal information loses its grip. We conclude, also in light of Figure 1, that the analysis of compatibility is not enough and it should be coupled with an analysis of the consistency of the preferences. Starting from the principle that the level of inconsistency of preferences should not be model-dependent, therefore, we adopt the Ordinal Consistency Ratio ($OR$) and the Cardinal Consistency Ratio ($CR$) proposed by Liang et al., (2020) to quantify deviations from these consistency conditions in the BWD method.
@@
3.4 Extreme ranking analysis
A number of sensitivity analysis tools already proposed for robust ordinal regression can be easily implemented for the BWD model. One example is the extreme ranking analysis Kadziński et al., (2012). In some decision-making contexts, it is often important to determine the maximum and minimum possible ranks that can be achieved by an alternative in the final ranking. This type of sensitivity analysis helps assess the stability of an alternative’s position. To determine these bounds, we solve two optimization problems for each alternative, always searching within the set of the optimal solutions of (BWD).
@@ 3.5 An interval-valued extension
subject to $-\xi+a_{Bi}^{-}V_{L}(\mathbf{x}_{i})\leq V_{L}(\mathbf{x}_{B})\leq a_{Bi}^{+}V_{L}(\mathbf{x}_{i})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
$-\xi+a_{iW}^{-}V_{L}(\mathbf{x}_{W})\leq V_{L}(\mathbf{x}_{i})\leq a_{iW}^{+}V_{L}(\mathbf{x}_{W})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
@@
Proposition 1.
If $a_{Bi}\in[a_{Bi}^{-},a_{Bi}^{+}]$, $a_{iW}\in[a_{iW}^{-},a_{iW}^{+}]$ and the consequence vectors are the same for both models, then, $\xi^{*}_{I}\leq\xi^{*}$.
Proof.
Let us consider the case of the comparisons between the best and the others. In the I-BWD, the first family of constraints is a linearization of
$\xi_{I}\geq\max\left\{V(x_{B})-a^{+}_{Bi}V(x_{i}),a^{-}_{Bi}V(x_{i})-V(x_{B}),0\right\}$ (28)
In the BWD, if we consider the real value $a_{Bi}$ as a degenerate interval, i.e. $a_{Bi}\in[a_{Bi},a_{Bi}]$, then we can write
$\xi\geq\max\left\{V(x_{B})-a_{Bi}V(x_{i}),a_{Bi}V(x_{i})-V(x_{B}),0\right\}.$ (29)
As we know that $a_{Bi}^{-}\leq a_{Bi}\leq a_{Bi}^{+}$ and $V(\mathbf{x}_{i})\geq 0$ for all $i\in R$ we deduce
$V(x_{B})-a^{+}_{Bi}V(x_{i})$ $\leq V(x_{B})-a_{Bi}V(x_{i})$
$a^{-}_{Bi}V(x_{i})-V(x_{B})$ $\leq a_{Bi}V(x_{i})-V(x_{B})$
and therefore the right-hand side of (29) cannot be smaller than the right-hand side of (28). We skip the formal proof, but a similar line of reasoning can be applied to the second family of constraints for both (BWD) and (I-BWD). If we consider that $\xi\geq 0$ naturally holds in (BWD), then we can see that all other constraints are the same, guaranteeing the same freedom in the definition of $V$. Hence, due to the minimization of both $\xi_{I}$ and $\xi$, we obtain $\xi_{I}^{*}\leq\xi^{*}$. ∎
@@ 4 A case study in logistics performance evaluation
We tested the BWD on the data collected and used by Rezaei et al., (2018) to rank countries according to a composite indicator of logistic performance. In their study, data from the World Bank was used to assess the performance of the World countries with respect to the following criteria: Customs; Infrastructure; Services; Timeliness; Tracking and tracing; International shipment. For the sake of simplicity and ease of interpretation of the results, we only focused on European countries.
@@
One splitting point was allowed for each criterion set and was positioned midway in the interval $[\underline{x}_{j},\overline{x}_{j}]$. Next, we solved the set covering-inspired optimization problem (15)–(18) to find a mutually non-dominated minimum covering set. This set contained only three countries and was taken as a good initial suggestion. Given the possibility of asking for more comparisons, we added two more alternatives to better explore the lower part of the ranking. The final reference set $R$ was defined as
@@
Using the Ordinal Consistency Ratio ($OR$) as formulated in eq. (20), we found $OR=0.2$, indicating a violation of ordinal consistency. The local Ordinal Consistency Ratios ($OR_{i}$), shown in Table 2, help locate specific inconsistencies. In this case, the preferences provided by the expert on Latvia are not ordinally consistent with those expressed on Greece.
@@
After identifying and locating inconsistencies in the preferences, we re-engaged with the expert to review and reconsider her judgments. Upon reflection, the expert acknowledged the inconsistencies and proceeded to modify the preferences using the consistency improvement process from the work of Liang et al., (2022). Using the same approach, we derived the admissible ranges for $A^{BO}$ and $A^{OW}$, as shown in Table 3, and the visualization of the improving ranges, shown in Figure 2.
@@
With the help of the visualization of the ranges for improvement, the expert can easily adjust the inconsistent comparison. After consideration, the expert made some minor revisions on the preferences. The modified sets $A^{BO}$ and $A^{OW}$ are reported in Table 4.
@@
The full analysis revealed that $\succ_{\mathcal{N}}$ is almost a total order relation. This is confirmed in Figure 3, showing very low variability in the set of rankings obtained through optimization problems (22) and (23). Only one swap is possible, between Greece and Slovenia. The results shown in Figure 3 seem to hint at a good precision of the outcome and further tests that we conducted, but we avoided reporting, showed any level of incompatibility, i.e. $\xi^{*}>0$, very often leads to an extremely stable ranking.
@@
We also tested the I-BWD on the same case study, together with the analysis of dominance relations and maximum and minimum ranks. In this case, the expert information is summarized in Table 5. Note that the improving ranges shown in Table 3 can be used to guide experts to extend their original real-valued preferences to interval ones in a way that may decrease their inconsistency.
@@
Results of the extreme ranking analysis are presented in the right-hand side of Figure 3 and confirms the previous intuition on the much greater variability of the attainable ranks. To quantify the imprecision due to the existence of multiple compatible rankings, we introduce the following heuristic index
@@
where each term of the sum is the “spread” of the possible ranks that the $i$th alternative can attain divided by the maximum possible range, going from $1$, the best rank, to $m$, the lowest possible attainable rank. The scaling factor preceding the sum normalizes the quantity into the interval $[0,1]$. In the case of the I-BWD, we have $U=0.05938$, whereas, for the real-valued BWD it was $U=0.00135$.
@@ 5 Conclusions
The Best-Worst Disaggregation Method (BWD) offers an advancement in the field of Multi-Criteria Decision Analysis (MCDA) by combining the principles of the Best-Worst Method (BWM) with the disaggregation approach. This method addresses the challenges inherent in traditional aggregation and disaggregation techniques by leveraging the strengths of both to create a robust and reliable decision-making framework. It is a valuable tool in scenarios where experts have a clear understanding of alternatives but may not have predefined value functions.
@@
Future research on BWD could adapt it to accommodate group decision-making scenarios by developing techniques to aggregate individual preferences and address potential conflicts; adapting BWD for use in dynamic environments where decision criteria and alternatives may change over time; enhancing the computational efficiency of BWD to handle larger sets of criteria and alternatives by exploring advanced optimization techniques and parallel computing strategies; investigating methods to minimize biases in expert judgments, such as incorporating techniques from behavioral decision theory or utilizing machine learning or artificial intelligence to identify and correct inconsistencies; conducting practical case studies in various domains to validate the effectiveness of BWD and identify specific areas for improvement. Finally, in this manuscript we employed the well-known 1–9 scale, but other methods to collect statements on alternatives, such as those used in GRIP (Figueira et al., , 2009), may be incorporated.